OpenStudy (xochital):

help with 4 expressions please?

OpenStudy (xochital):

first one: \[\frac{ x }{ x ^{2} - 2 } = \frac{ 2 }{ 3x + 4 }\] what is the value of x?

OpenStudy (anonymous):

x(3x+4)=2(x^2-2)

OpenStudy (anonymous):

3x^2+4x=2x^2-4

OpenStudy (mathstudent55):

For the first expression, cross multiply.

OpenStudy (xochital):

so would i combine like terms now?

OpenStudy (mathstudent55):

Bring all terms to left side. Combine them and solve the quadratic equation.

OpenStudy (xochital):

so would it be \[3x ^{2} + 4x +2x ^{2} -4 \] ?

OpenStudy (mathstudent55):

\(3x^2+4x=2x^2-4 \) \(3x^2 - 2x^2 + 4x + 4 = 0\)

OpenStudy (xochital):

is that how i can find the value of x?